McGonigal’s Meats, Inc. currently pays no dividends. The firm plans to begin paying dividends in 3 years (at the end of t3). The first dividend at that time will be $1 and dividends are expected to grow at 5% per annum thereafter.

Given shareholders demand a 12% return on their investments, what is the price of the

stock today (t0)?
1
Expert's answer
2014-04-24T11:00:59-0400
Generally, investors buy common stocks for two reasons. Many stocks offer a cash dividend, and they also have the potential to provide a capital gain. In this publication, we will present a method for calculating stock prices based on a constant growth model, leveraging a discounted cash flows approach which considers both dividends and capital gains.
Expected Return = (Dividends Paid + Capital Gain) / Price of Stock
This expected return for a stock is also known as the market capitalization rate or discount rate. We're going to use all three terms interchangeably throughout our calculations and explanations in this article. Let's look at a quick example of how this formula works.
P(t3) = 1(1.05)/(0.12-0.05) = 15
P(t0)= (15+4)/1.12^3 = $13.52